OpenStudy (anonymous):

Use Laplace transformations to solve the IVP: x''+6x'+9x=7e^(4t) x(0)=0, x'(0)=0

OpenStudy (amistre64):

well, we simply need to get our hands dirty and take the Laplace to start with

OpenStudy (amistre64):

ideally, we should have built up a table of laplace transforms to be able to speed up the process.

OpenStudy (sidsiddhartha):

yes agree with you

OpenStudy (sidsiddhartha):

\[L[\frac{ dy }{ dt }]=sY(s)-Y(0)\]

OpenStudy (amistre64):

\[\int_{0}^{inf} fe^{-st}~dt\] making a table of by parts e^(-st) + f -e^(-st)/s - f' e^(-st)/s^2 + f'' -e^(-st)/s^3 - f''' e^(-st)/s^4 .... we can develop the general setup

OpenStudy (anonymous):

After doing the transformation, this is what i got and am not sure about it \[y=\frac{ s^2-4s+7 }{ (s-4)(s+3)^2 }\]

OpenStudy (sidsiddhartha):

\[L[Y'']=s^2Y(s)-sY(0)-Y'(0)\]

OpenStudy (sidsiddhartha):

now lets apply LT to the equation

OpenStudy (amistre64):

yes, the inverse :)

OpenStudy (sidsiddhartha):

\[x''+6x'+9x=7e^{4t}\\apply~LT\\s^2X(s)-sX(s)-X'(0)+6sX(s)-X(0)+9X(s)=\frac{ 7 }{ s-4 }\]

OpenStudy (sidsiddhartha):

seems ok?

OpenStudy (sidsiddhartha):

now put the initial conditions

OpenStudy (anonymous):

does that look okay \[(s^2+6s+9)X=\frac{ 7 }{ s-4 }\]

OpenStudy (anonymous):

\[X=\frac{ 7 }{ ((s-4)(s+3)^2 }\]

OpenStudy (sidsiddhartha):

\[s^2X(s)-sX(s)+6sX(s)+9X(s)=\frac{ 7 }{ s-4 }\\X(s)(s^2+5s+9)=\frac{ 7 }{ s-4 }\]

OpenStudy (sidsiddhartha):

u missed a s it should be 5s not 6s okay?

OpenStudy (anonymous):

Doesnt that s cancel since we have x(0)=0 for the initial value.

OpenStudy (sidsiddhartha):

oh yes sorry my bad ,

OpenStudy (sidsiddhartha):

so u have now\[X(t)=L^{-1}\frac{ 7 }{( s-4)(s+3)^2 }\] now we have to use partial fraction

OpenStudy (anonymous):

Then how do i get A, B and C \[\frac{ A}{ s-4 }+\frac{ B }{ s+3}+\frac{ C }{ (s+3)^2 }\]

OpenStudy (sidsiddhartha):

or we can use convolution ,are u familiar with it?

OpenStudy (anonymous):

I havent heard of that.

OpenStudy (sidsiddhartha):

ok lets do partial fraction ,ure attemp is correct

OpenStudy (sidsiddhartha):

ok can u do it ? the expansion?

OpenStudy (anonymous):

So far av been able to get A=1. The rest am confused how to go about it

OpenStudy (anonymous):

S

OpenStudy (sidsiddhartha):

\[\frac{ 7 }{ (s-4)(s+3)^2}=\frac{ A}{ s-4 } +\frac{ B }{ s+3 }+\frac{ c }{( s+3)^2 }\\ \frac{ A.(s+3)^2+B.(s-4)(s+3)+C(s-4) }{ (s+3)^2.(s-4) }=\frac{ 7 }{ (s-4)(s+3)^2 }\\A(s+3)^2+b(s+3)(s-4)+C(s-4)=7\\now ~put~,s=-3\\we~get\\C(-3-4)=7\\C=-1\]

OpenStudy (sidsiddhartha):

now wait a bit for A,B

OpenStudy (sidsiddhartha):

now \[put~s=4~we'\ll ~have\\A(3+4)^2=7\\A=1/7\]

OpenStudy (sidsiddhartha):

now forB\[put~s=1\\A(1+3)^2+B(1+3)(1-4)+C(1-4)=7\\ \frac{ 16 }{ 7 }-12B+3=7\\B=\frac{ -1 }{ 7 }\]

OpenStudy (sidsiddhartha):

\[so~We~have\\A=1/7\\B=-1/7\\C=-1\]

OpenStudy (sidsiddhartha):

ok????

OpenStudy (sidsiddhartha):

now i believe u can do the rest

OpenStudy (anonymous):

So far it looks okay. I had made an algebra error on A but i got it now

OpenStudy (sidsiddhartha):

ok u can take it from here?

OpenStudy (anonymous):

Yes and now i can be able to do the rest of the homework. Thanks alot

OpenStudy (sidsiddhartha):

no problem buddy :)